# Digital Histopathology with Graph Neural Networks: Concepts and Explanations for Clinicians

### 04 Dec 2023 v1

@@ 4 Discussion
These explainability methodologies afford engineers, in collaboration with clinicians, the ability to fine tune parameters. On the one hand, a number of concepts $k$ can lead to higher concept completeness scores [19], which is a measure for how powerful prediction is using the concepts alone. This comes at the cost of: $(i)$ decreasing quality of clustering as observed by a reduction in average silhouette scores [20] from $0.355$ at $k=4$ to $0.191$ at $h=512$, $(ii)$ a more complicated and longer instance-level explanations due more min-terms. To some extent, using a harsher $L_{1}$ regularization can reduce explanation complexity, but this only throws away information in the graph-level concept and we poorer accuracy was expected.

### 28 Dec 2023 v2

@@ 4 Discussion
These explainability methodologies afford engineers, in collaboration with clinicians, the ability to fine tune parameters. On the one hand, a number of concepts $k$ can lead to higher concept completeness scores [19], which is a measure for how predictive power using the concepts alone. This comes at the cost of: $(i)$ decreasing quality of clustering as observed by a reduction in average silhouette scores [20] from $0.355$ at $k=4$ to $0.191$ at $h=512$, $(ii)$ a more complicated and longer instance-level explanations due more min-terms. We found that increasing the degree of $L_{1}$ regularization assisted in both improving accuracy as well as reducing explanation complexity.